# Review of Degenerate Higher Order Scalar Tensor Theories in Cosmology

### v1

@@ Abstract
Degenerate Higher Order Scalar Tensor (DHOST) theories are the most general scalar-tensor theories whose Lagrangian depends on the metric tensor and a single scalar field and which propagate only one scalar degree of freedom, without being plagued by Ostrogradsky instabilities. This is achieved through certain degeneracies of the functions forming their Lagrangian. They generalise the Horndeski and beyond-Horndeski theories. Originally build to describe the late-time acceleration of the expansion of the universe, generalising the cosmological constant, they can be used to build models for early universe, to describe inflation or alternatives to standard inflation. In the late universe, they modify the standard Vainstein screening mechanism from Horndeski theories which can have observable consequences and are suited to build black hole models, featuring non-stealth Kerr black hole solutions. In this work we review their effects on cosmology, looking at their basic properties, their parameterisations and classifications, and then focusing on solutions in the early and the late universe.
@@ 1 Introduction
The General Theory of Relativity, proposed more than a century ago, supplemented by a cosmological constant, has been successfully used to describe the universe. Its predictions are in agreement with the observations, including the accelerated expansion of the universe and the presence of gravitational waves.
The nature of the cosmological constant is, however, unexplained and this has led to a large number of theories being developed, trying to relate this to the unknown nature of the dark energy, but also to find a way to parameterise it and to constrain it with various cosmological probes. Moreover, extensions of the theory (modified gravity) are required to account for the presence of singularities such as black holes.
In the very early universe, just after in the initial Big bang, there was a brief period of exponential expansion, the inflationary epoch, developed from the late 1970s by several theoretical physicists including Alexei Starobinsky, Alan Guth, Stephen Hawking and Andrei Linde. This theory provides a satisfactory explanation for the presence of the structures that we observe in the universe today through the rapid growth of primordial quantum fluctuations [1, 2, 3, 4].
As these two early and late universe epochs are qualitatively similar, theories that are developed for one era can be adapted to the other. A classical method to parameterise such theories is to introduce scalar and tensor fields to model the interactions. In this work we concentrate on theories involving one single scalar field. It was long thought that the most general such theories that do not yield additional degrees of freedom and hence avoid the generation of Ostrogradsky ghost instabilities [5] need to have second-order Euler-Lagrange equations. These are the Horndeski theories [6]. It turns out that this last requirement is not necessary, and they have been generalised to beyond Horndeski [7, 8, 9] and finally to the Degenerate Higher Order Scalar Tensor (DHOST) theories [10]. The latter include all the scalar-tensor theories having these properties, including Horndeski and beyond Horndeski ones.
@@
The early developments of DHOST theories are reviewed in [11], where the consequences of these theories in the late universe and on dark energy are discussed. The paper presents the emergence of DHOST theories through the generalisation of Horndeski and beyond Horndeski ones and discusses their cosmological consequences as well as the gravitational properties of astrophysical bodies in this context. The mechanisms concerning the degeneracies are presented in detail, followed by a classification of the models and then the study of the theories within the formalism of the Effective Field Theory of dark energy and screening mechanisms.
@@
In this review, we discuss the uses of DHOST theories in cosmology in both phases of accelerated expansion of the universe. We introduce the DHOST theories in Section 2, and briefly mention the traditional classification of quadratic DHOST theories. The Effective Field Theory of Dark Energy and its implications on DHOST models is then presented in Section 3. The consequences of DHOST theories on astrophysical bodies (neutron stars and black holes) and the Vainshtein mechanism are the subject of Section 4. The DHOST model for inflation is presented in Section 5, while bouncing cosmology with DHOST models is reviewed in Section 6. We conclude in Section 7.
@@ 2 DHOST Theories
DHOST theories involve derivatives up to second order in the scalar field $\phi$ and can be classified in terms of the powers of derivatives present. So far, quadratic [10, 12, 13, 14] and cubic [15] theories have been studied. The most general action containing up to cubic terms is given by
@@
$\phi_{\mu}\equiv\nabla_{\mu}\phi\,,\qquad\phi_{\mu\nu}\equiv\nabla_{\nu}\nabla_{\mu}\phi\,,\qquad X\equiv\phi^{\mu}\phi_{\mu}\,$ (2.2)
@@
We note that Horndeski theories are obtained by setting
$F_{(2)}=G_{4}\,,\qquad A_{1}=-A_{2}=2G{4,X}\,,\qquad A_{3}=A_{4}=A_{5}=0\,,$
@@
As these results are very technical, I shall present in the following only the classification of quadratic theories, leaving the interested reader to discover the classification of the cubic and mixed quadratic/cubic theories in [15].
@@
Here we have dropped the subscript from the function $F_{(2)}$, as we are only dealing with quadratic theories. The classification of the theories is given in terms of how these three equations can be satisfied simultaneously. We express in square brackets the naming used in Ref. [12]
@@ 3 Effective Field Theory of Dark Energy and cosmological constraints
Parameters $M$, $\alpha_{T}$, $\alpha_{K}$, $\alpha_{B}$ are specific to Horndeski theories, while the additional parameter $\alpha_{H}$ appears in beyond-Horndeski theories. The physical description of these parameters is discussed in [28].
@@
For the scalar, that is
@@
In the case of $\mathcal{C}_{\rm II}$ theories ${B}_{\tilde{\zeta}}=-2(1+\alpha_{T})$, which combined with the result for the tensor modes, shows that scalar and tensor modes cannot be stable simultaneously, and hence this class of theories is not viable.
@@
This restriction is based on the assumption that scales probed by the LIGO-Virgo measurements (which are of the order of $10^{3}$ km), can be extrapolated by several orders of magnitude to cosmological scales.
@@
The EFT method has been used, together with the CMB likelihoods from Planck, to place constraints on the EFT parameters in DHOST models [38, 39].

### v3

# Recent Developments in Degenerate Higher Order Scalar Tensor Theories

@@ Abstract
Degenerate Higher Order Scalar Tensor (DHOST) theories are the most general scalar-tensor theories whose Lagrangian depends on the metric tensor and a single scalar field and its derivatives up to second order. They propagate only one scalar degree of freedom, without being plagued by Ostrogradsky instabilities. This is achieved through certain degeneracies of the functions forming their Lagrangian. They generalise the Horndeski and beyond-Horndeski theories. Originally proposed to describe the late-time acceleration of the expansion of the universe, generalising the cosmological constant, they can also be used to build models of the early universe, to describe inflation or alternatives to standard inflation. In the late universe, they modify the standard Vainstein screening mechanism from Horndeski theories (which can have observable consequences) and are suited to build black hole models, featuring non-stealth Kerr black hole solutions. In this work their phenomenology is reviewed, looking at their basic properties, their parameterisations and classifications, focusing on solutions in the early and the late universe and at cosmological and astrophysical constraints.
@@ 1 Introduction
The General Theory of Relativity (GR), proposed more than a century ago, supplemented by a cosmological constant, has been successfully used to describe the universe. Its predictions are in agreement with observations, including the accelerated expansion of the universe [1, 2] and the presence of gravitational waves [3].
The standard cosmological model, $\Lambda$CDM, consists of dark energy in the form of a cosmological constant ($\Lambda$), cold dark matter and ordinary matter. The cosmological constant, originally proposed by Einstein, leads to the so-called “cosmological constant problem”, a discrepancy of 120 orders of magnitude between the value of the vacuum energy predicted by quantum field theory and the observed one from cosmology [4]. In strong gravitational fields, such as in the vicinity of black hole singularities, predictions of GR are no longer accurate. These issues have led to a large number of alternative theories being developed, trying to relate the expansion of the universe to dark energy, which is also of an unknown nature, and also to find a way to parameterise it and to constrain it with various cosmological and astrophysical probes. The main alternatives are models of dark energy [5] and modified gravity [6]. Lovelock’s theorem [7, 8] states that GR is the only 4D theory that has second-order Euler-Lagrange equations that can be build only from the metric. Hence, extensions or alternatives to GR must involve at least an additional degree of freedom. Various such models have been developed, including models depending on a single additional scalar field: Brans-Dicke [9], quintessence [10], kinetic dominated scalar fields (k-essence) [11, 12, 13], chameleons [14], dilatons [15], galileons [16], covariant galileons [17]. Other models have also been studied, see e.g. Refs. [18, 19] for reviews and the classification of models of modified gravity. A more general class of models is represented by Horndeski theories [20, 21], of which the previous mentioned theories are special cases. These theories are characterised by second-order Euler-Lagrange equations, which was thought to be a requirement for the theories to be ghost-free (in order for the theory to be stable, one needs to avoid the generations of Ostrogradsky ghost instabilities [22] which yield additional degrees of freedom). Therefore, much of the interest has focused on theories which have second-order equations of motion. It turns out that this last requirement is not necessary, and thus they have been generalised to beyond Horndeski (Gleyzes-Langlois-Piazza-Vernizzi) [23, 24, 25] and finally to the Degenerate Higher Order Scalar Tensor (DHOST) theories, first introduced in [26]. These are the most general scalar-tensor theories that propagate one scalar degree of freedom and which involve up to second-order derivatives of a single scalar field in their Lagrangians. They include all the scalar-tensor theories having these properties, including the Horndeski and beyond Horndeski ones.
In parallel, the origin and the first moments of the universe are also of great interest. Similarly to the late-time epoch previously described, the universe also underwent a brief period of accelerated expansion, known as inflation, just after the initial Big Bang. The theory describing it was developed from the late 1970s by several theoretical physicists including Alexei Starobinsky, Alan Guth, Stephen Hawking and Andrei Linde. The theory was originally built to solve various cosmological puzzles, such as the horizon problem, the flatness problem and the magnetic monopole problem. It has subsequently been shown that it can also be used to explain the presence of the structures that we observe in the universe today through the rapid growth of primordial quantum fluctuations [27, 28, 29, 30]. This epoch has been constrained by cosmological probes such as Planck [31]. Its nature and the mechanisms which govern it remain nevertheless unknown and are active fields of research [32]. Many of the models describing inflation are based on the presence of at least a scalar field, the inflaton. Additionally, classical inflation usually assumes a ‘Big Bang’ moment at the beginning of the universe, associated with a singularity. This assumption is relaxed in various alternative theories to inflation [33, 34, 35]. The similar characteristics of the two epochs – inflation and late-time accelerated expansion of the universe – make the use of similar types of models appropriate and hence theories that are developed for one era can be adapted to the other [18].
@@
The early developments of DHOST theories are reviewed in [36]. The paper presents the emergence of DHOST theories through the generalisation of Horndeski and beyond Horndeski ones and discusses their cosmological consequences as well as the gravitational properties of astrophysical bodies in this context. The mechanisms concerning the degeneracies are presented in detail, followed by a classification of the models and then the study of the theories within the formalism of the Effective Field Theory of dark energy and screening mechanisms. Results related to DHOST theories in the late universe and black hole solutions are also reviewed in [37].
@@
In this review, we discuss the phenomenological consequences of DHOST theories in cosmology in both phases of accelerated expansion of the universe, and also for astrophysical objects. We introduce the DHOST theories in Section 2, and briefly mention the traditional classification of the quadratic ones. The Effective Field Theory of Dark Energy and its implications on DHOST models is then presented in Section 3. The consequences of these theories on astrophysical bodies (neutron stars and black holes) and the Vainshtein mechanism are the subject of Section 4. Other aspects, both cosmological and astrophysical, as presented in the literature, are mentioned in Section 5. A DHOST model for inflation is presented in Section 6, while an alternative model, DHOST Genesis, is reviewed in Section 7. We conclude in Section 8.
@@ 2 DHOST Theories
DHOST theories involve derivatives up to second order in a scalar field $\phi$ and can be classified in terms of the powers of derivatives present. So far, quadratic [26, 38, 39, 40] and cubic [41] theories have been studied. Systematically classifying and analysing theories beyond third order becomes complicated due to the large number of terms involved. However, extensions to such theories have been studied phenomenologically. In Refs. [42, 43] a generalisation of the so-called mimetic gravity model has been discussed. More recently in Refs. [44, 45, 46] another extension has been proposed by performing a higher-derivative generalisation of disformal transformations on the Horndeski action. The most general action containing up to cubic terms is given by
@@
$\phi_{\mu}\equiv\nabla_{\mu}\phi\,,\qquad\phi_{\mu\nu}\equiv\nabla_{\nu}\nabla_{\mu}\phi\,,\qquad X\equiv\phi^{\mu}\phi_{\mu}\,,$ (2.2)
@@
We note that the Horndeski theories are obtained by setting
$F_{(2)}=G_{4}\,,\qquad A_{1}=-A_{2}=2G_{4,X}\,,\qquad A_{3}=A_{4}=A_{5}=0\,,$
@@
As these results are very technical, we will present in the following only the classification of quadratic theories, leaving the interested reader to discover the classification of the cubic and mixed quadratic/cubic theories in [41].
@@
Here we have dropped the subscript from the function $F_{(2)}$, as we are only dealing with quadratic theories. The classification of the theories is given in terms of how these three equations can be satisfied simultaneously. We express in square brackets the naming used in Ref. [38].
@@ 3 Effective Field Theory of Dark Energy and Cosmological Constraints
Parameters $M$, $\alpha_{T}$, $\alpha_{K}$, $\alpha_{B}$ are specific to Horndeski theories, while the additional parameter $\alpha_{H}$ appears in beyond-Horndeski ones. The physical description of these parameters is discussed in [59]. For quadratic and cubic theories, these EFT parameters can be expressed in terms of the $A_{i}$ and $B_{i}$ parameters as [54]
${M^{2}}=$ $\ 2F_{2}-X[2A_{1}+F_{3\phi}-6(B_{2}+B_{3})H\mu^{2}]\,,$ (3.8)
$\alpha_{T}=$ $-1+\frac{2F_{2}+XF_{3\phi}}{M^{2}}\;,$ (3.9)
$\alpha_{H}=$ $-1+\frac{1}{M^{2}}\left[2F_{2}-X\left(4F_{2X}+2F_{3X}H\mu^{2}+F_{3\phi}\right)\right]\,,$ (3.10)
$\alpha_{L}=$ $\ \frac{3X}{M^{2}}\left[A_{1}+A_{2}-(9B_{1}+5B_{2}+3B_{3})H\mu^{2}\right]\;,$ (3.11)
$\beta_{1}=$ $\ \frac{X}{2M^{2}}\left\{4F_{2X}+2A_{2}+XA_{3}+2\left[2F_{3X}-9B_{1}-B_{2}-X(3B_{4}+B_{6})\right]H\mu^{2}\right\}\,,$ (3.12)
$\beta_{2}=$ $-\frac{2X}{M^{2}}\left\{A_{1}+A_{2}+X(A_{3}+A_{4})+X^{2}A_{5}-3\left[3B_{1}+B_{2}+X(2B_{4}+B_{5})+X^{2}B_{9})\right]H\mu^{2}\right\}\,,$ (3.13)
$\beta_{3}=$ $-\frac{2X}{M^{2}}\left\{4F_{2X}-2A_{1}-XA_{4}+\left[4F_{3X}+6B_{2}+3B_{3}+X(3B_{5}+B_{7})\right]H\mu^{2}\right\}\,,$ (3.14)
where a linear assumption $\phi=\mu^{2}t$ is made.
@@
In particular, from Eqs. (3.8) and (3.9) one can see that the condition $\alpha_{T}=0$ corresponds to $A_{1}=0$. For the scalar, the quadratic action becomes
@@
In the case of $\mathcal{C}_{\rm II}$ theories, ${B}_{\tilde{\zeta}}=-2(1+\alpha_{T})$, which combined with the result for the tensor modes, shows that scalar and tensor modes cannot be stable simultaneously, and hence this class of theories is not viable.
@@
This restriction is based on the assumption that scales probed by the LIGO-Virgo measurements (which are of the order of $10^{3}$ km), can be extrapolated by several orders of magnitude to cosmological scales. We note that, if the decay rate is too large, we would not have observed gravitational waves. In Ref. [64], the authors found that imposing $A_{3}=0$ on top of $A_{1}=0$ leads to a suppression of the decay rate of gravitational waves such that the decay time can be longer than the age of the Universe.
@@
We illustrate in this section the cosmological evolution of quadratic DHOST theories, as described in Ref. [69]. In this work, the gravitational wave speed constraints are not used, as no extrapolation from astrophysical to cosmological scales is assumed. The aim is to find self-accelerating solutions, corresponding to the late-time evolution of the universe and the transition between the matter and accelerated expansion epochs.
The starting point is a homogeneous background, depending on a lapse function $N$ and a scale factor $a$,
$ds^{2}=-N^{2}(t)dt^{2}+a^{2}(t)\delta_{ij}dx^{j}dx^{i}\,,$ (3.24)
The homogeneous action becomes:
$S_{\rm hom}[N,a,\phi]$ $=$ $\int dt\,a^{3}N\Biggl{\{}P+Q\left(\frac{\dot{N}}{N^{3}}\dot{\phi}-\frac{3\dot{a}}{a\,N^{2}}\dot{\phi}-\frac{\ddot{\phi}}{N^{2}}\right)-F_{\phi}\frac{6\dot{a}}{a\,N^{2}}\dot{\phi}$ (3.25)
$\qquad\qquad-\frac{6f_{1}}{N^{2}}\left(\frac{\dot{a}}{a}+\frac{f_{2}}{4f_{1}}\left(\frac{\dot{N}\dot{\phi}{}^{2}}{N^{3}}-\frac{\ddot{\phi}\dot{\phi}}{N^{2}}\right)\right)^{2}\Biggr{\}}\,,$
where
$f_{1}\equiv F-XA_{1}\,,\qquad f_{2}\equiv 4F_{X}-2A_{1}+XA_{3}\,,$ (3.26)
and
$X=-\frac{\dot{\phi}{}^{2}}{N^{2}}\,.$ (3.27)
The equations of motion from (3.25) appear to be higher than second order, but this is not the case due to the degeneracy conditions. This can be seen by defining an auxiliary scale factor $b$,
$a\equiv\Lambda\left(X,\phi\right)b\equiv e^{\lambda(X,\phi)}b\,.$ (3.28)
Here $\lambda$ satisfies
$\lambda_{X}=-\frac{f_{2}}{8f_{1}}\,,$ (3.29)
and hence the quadratic terms in $\ddot{\phi}$ in the action (3.25) are reabsorbed in the derivatives of the new scale factor.
The Hubble parameter associated with this auxiliary scale factor is defined as
$H_{b}\equiv\frac{\dot{b}}{Nb}=H-\lambda_{X}\frac{\dot{X}}{N}-\lambda_{\phi}\frac{\dot{\phi}}{N}\,,\qquad\dot{X}=\frac{2}{N^{2}}\left(\frac{\dot{N}\dot{\phi}^{2}}{N}-{\dot{\phi}\ddot{\phi}}\right)\,.$ (3.30)
From action, (3.25) one can express the Lagrangian in terms of the new expansion factor $b$ as
$L_{\rm hom}$ $=$ $\Lambda^{3}b^{3}N\Biggl{\{}-3\lambda_{\phi}\frac{\dot{\phi}{}^{2}}{N^{2}}(2f_{1}\lambda_{\phi}+Q+2F_{\phi})+P-6f_{1}H_{b}^{2}-\frac{3\dot{\phi}}{N}\left(4\lambda_{\phi}f_{1}+2F_{\phi}+Q\right)H_{b}$ (3.31)
$\qquad+\left(Q-6\lambda_{X}\left(2F_{\phi}+Q\right)\frac{\dot{\phi}2}{N^{2}}\right)\left(\frac{\dot{N}\dot{\phi}}{N^{3}}-\frac{\ddot{\phi}}{N^{2}}\right)\Biggr{\}}.$
The total Lagrangian $L$ is then the sum of the preceding one and the matter Lagrangian $L_{m}$. The equations of motion are obtained by considering the Euler-Lagrange equations for $N$, $b$ and $\phi$. They yield generalised Friedmann equations and the scalar field equation of motion respectively,
$-\frac{d^{2}}{dt^{2}}\frac{\partial L}{\partial\ddot{\phi}}+\frac{d}{dt}\frac{\partial L}{\partial\dot{\phi}}-\frac{\partial L}{\partial\phi}$ $=$ $0$ (3.32)
Then one can fix the time coordinate such that $N=1$ (and $\dot{N}=0$).
The matter is assumed to be described by a perfect fluid whose equation of state is $P=w\rho$, where $w$ is constant. The energy-momentum tensor of the fluid is defined as
$T^{\mu\nu}=\frac{2}{\sqrt{-g}}\frac{\delta S_{m}}{\delta g_{\mu\nu}}\,,\qquad S_{m}=\int d^{4}x\sqrt{-g}\,L_{m}\,.$ (3.33)
Then the variation of the matter Lagrangian is given by
$\delta L_{m}=-a^{3}\rho_{m}\delta N+3Na^{2}P_{m}\delta a\,,$ (3.34)
where $\rho_{m}$ and $P_{m}$ are the fluid energy density and pressure, respectively.
The two Friedmann equations take the form
$g_{0}+g_{1}H_{b}\dot{\varphi}+g_{2}H_{b}^{2}$ $=\left(1-6w\lambda_{X}\dot{\varphi}^{2}\right)\rho_{m}\,,$ (3.35)
$g_{3}+g_{4}(2\dot{H}_{b}+3H_{b}^{2})+g_{5}H_{b}\dot{\varphi}+g_{6}\ddot{\varphi}+g_{7}H_{b}\dot{\varphi}\ddot{\varphi}$ $=-w\rho_{m}\,.$ (3.36)
and the scalar field equation can be written as
$\frac{d}{dt}\left(b^{3}\Lambda^{3}J\right)+b^{3}\Lambda^{3}U=0\,,$ (3.37)
where
$U=g_{8}+g_{9}H_{b}\dot{\varphi}+g_{10}H_{b}^{2}+g_{11}\ddot{\varphi}\,,\qquad J=g_{12}\dot{\varphi}+g_{13}H_{b}+g_{14}H_{b}^{2}\dot{\varphi}\,.$ (3.38)
The coefficients $g_{i}$ can be expressed in terms of quantities in the Lagrangian and of $\lambda$ and are given in Appendix A of Ref. [69].
To investigate the stability of the solutions of these equations, one needs to consider linear perturbations, and this has been described in the previous paragraphs. In the absence of matter, the stability conditions are given in (3.22). If one also considers matter, then the analysis is significantly more complicated because there is an additional degree of freedom, $\delta\sigma$. Then the quadratic action can be expressed as
[…]
where the vector $V^{T}=(\zeta,H\frac{\delta\sigma}{\dot{\sigma}})$ contains the two scalar degrees of freedom and the dots stand for the terms with fewer than two (space or time) derivatives, which are not relevant for the stability discussion. The kinetic and gradient matrices read [66]
${\bf K}$ $=\left(\begin{matrix}A_{\zeta}+\frac{\rho_{m}(1+w_{m})}{M^{2}c_{m}^{2}\left(H(1+\alpha_{B})-\dot{\beta}_{1}\right)^{2}}\qquad&\frac{\rho_{m}(1+w_{m})\left(3c_{m}^{2}\beta_{1}-1\right)}{M^{2}c_{m}^{2}H\left(H(1+\alpha_{B})-\dot{\beta}_{1}\right)}\\[17.22217pt] \frac{\rho_{m}(1+w_{m})\left(3c_{m}^{2}\beta_{1}-1\right)}{M^{2}c_{m}^{2}H\left(H(1+\alpha_{B})-\dot{\beta}_{1}\right)}&\frac{\rho_{m}(1+w_{m})}{M^{2}c_{m}^{2}H^{2}}\end{matrix}\right)\,,$ (3.40)
${\bf G}$ $=\left(\begin{matrix}B_{\zeta}&-\frac{\rho_{m}(1+w_{m})\left(1+\alpha_{H}+(1+\alpha_{T})\beta_{1}\right)}{M^{2}H\left(H(1+\alpha_{B})-\dot{\beta}_{1}\right)}\\[17.22217pt] -\frac{\rho_{m}(1+w_{m})\left(1+\alpha_{H}+(1+\alpha_{T})\beta_{1}\right)}{M^{2}H\left(H(1+\alpha_{B})-\dot{\beta}_{1}\right)}&\frac{\rho_{m}(1+w_{m})}{M^{2}H^{2}}\end{matrix}\right)\,.$ (3.41)
To avoid ghost and gradient instabilities, both matrices $\bf K$ and $\bf G$ must be positive definite. If the matter satisfies $c_{m}\ll 1$ and $w_{m}\ll 1$, one can expand the expressions of the eigenvalues of $\bf K$ and $\bf G$ with respect to $c_{m}$ and $w_{m}$. At leading order, the eigenvalues of $\bf K$ and $\bf G$, $\lambda_{K_{1,2}}$ and $\lambda_{G_{\pm}}$ respectively, satisfy,
$\begin{split}\lambda_{K_{1}}&=\frac{A_{\zeta}M^{2}H^{2}(1+\alpha_{B}-\beta_{1}^{\prime})^{2}+6\rho_{m}\beta_{1}}{M^{2}H^{2}(1+(1+\alpha_{B}-\beta_{1}^{\prime})^{2})}\,,\quad\;\lambda_{K_{2}}=\frac{\rho_{m}}{c_{m}^{2}H^{2}M^{2}}\left[\frac{1}{(1+\alpha_{B}-\beta_{1}^{\prime})^{2}}+1\right],\\ \lambda_{G_{\pm}}&=\frac{B_{\zeta}}{2}+\frac{1}{2M^{2}}\left[\frac{\rho_{m}}{H^{2}}\pm\sqrt{\frac{4\rho_{m}^{2}(1+\alpha_{H}+(1+\alpha_{T})\beta_{1})^{2}}{H^{4}(1+\alpha_{B}-\beta_{1}^{\prime})^{2}}+(\frac{\rho_{m}}{H^{2}}-M^{2}B_{\zeta})^{2}}\,\right]\,.\end{split}$ (3.42)
Hence, $\lambda_{K_{2}}$ is always positive while the signs of the three other eigenvalues depend on the specific background solution.
In [69], specific numerical examples are investigated, explicitly showing that there are stable DHOST theories.
Another possible instability in such theories is the Dolgov-Kawasaki instability [70], which arises when a gravity theory allows for a dynamical equation for the scalar curvature through the trace of the metric field equations and the associated square mass is non-positive. In the case of Horndeski this has been shown not to be an issue [71].
The EFT method has been used, together with the CMB likelihoods from Planck, to place constraints on the EFT parameters in DHOST models [72, 73]. This work has relied upon the development of a Boltzmann solver that incorporates the EFT parameters and the use of a Markov-Chain Monte-Carlo method. Given the nature of the work, particular subclasses of DHOST have been chosen. For a simple model parameterised in terms of the dark energy density $\Omega_{\rm DE}$ as $\alpha_{\rm K}=\Omega_{\rm DE}(t)/\Omega_{\rm DE}(t_{0})$, $\alpha_{\rm B}=\alpha_{\rm T}=\alpha_{\rm M}=\alpha_{\rm H}=0$ and $\beta_{1}=\beta_{1,0}\Omega_{\rm DE}(t)/\Omega_{\rm DE}(t_{0})$ and a $\Lambda$CDM background, a constraint $\beta_{1,0}=0.032_{-0.016}^{+0.013}$ (at 68% CL) has been derived. For a theory given by $\mathcal{L}_{\rm DHOST}=X+c_{3}X\Box\phi/\Lambda^{3}+(M_{\rm Pl}^{2}/2+c_{4}X^{2}/\Lambda^{6})R+48c_{4}^{2}X^{2}/(M_{\rm Pl}^{2}\Lambda^{12}+2c_{4}\Lambda^{6}X^{2})\phi^{\mu}\phi_{\mu\rho}\phi^{\rho\nu}\phi_{\nu}$ with two positive constant parameters, $c_{3}$ and $c_{4}$, the authors have obtained $c_{3}=1.59^{+0.26}_{-0.28}$ and the upper bound on $c_{4}$, $c_{4}<0.0088$ (at 68% CL).